### v1

# Narrowing the Focus: Learned Optimizers for Pretrained Models

Abstract
In modern deep learning, the models are learned by applying gradient updates using an optimizer, which transforms the updates based on various statistics. Optimizers are often hand-designed and tuning their hyperparameters is a big part of the training process. Learned optimizers have shown some initial promise, but are generally unsuccessful as a general optimization mechanism applicable to every problem. In this work we explore a different direction: instead of learning general optimizers, we instead specialize them to a specific training environment. We propose a novel optimizer technique that learns a layer-specific linear combination of update directions provided by a set of base optimizers, effectively adapting its strategy to the specific model and dataset. When evaluated on image classification tasks, this specialized optimizer significantly outperforms both traditional off-the-shelf methods such as Adam, as well as existing general learned optimizers. Moreover, it demonstrates robust generalization with respect to model initialization, evaluating on unseen datasets, and training durations beyond its meta-training horizon.
1 Introduction
The optimization loop forms the backbone of machine learning algorithms. The choice of optimizer and its hyperparameters heavily influences the final model’s performance. Despite its importance, optimizer selection often relies on heuristics and domain-specific knowledge. For instance, Adam optimizer (Kingma & Ba, 2014) excels with language problems, but struggles with image classification problems. The potential of second-order methods for stochastic data remains largely untapped, except for a few notable exceptions (Martens & Grosse, 2015; Gupta et al., 2018). Every year, more and more optimization methods are proposed, making it increasingly challenging for practitioner to select the best one (Choi et al., 2019; Schmidt et al., 2021).
Learned optimizers or learning-to-learn methods (Schmidhuber, 1987; Bengio et al., 1990, 2013) offer a potential solution to this challenge by automating the optimizer selection process. These methods automate this process by either learning the hyperparameters of existing optimizers or developing entirely new optimization algorithms. However, two significant challenges limit widespread adoption of learned optimizers: meta-optimization difficulties, and meta-generalization.
Meta-optimization difficulties arise from the sensitivity of gradients in bi-level optimization with a high number of training iterations (aka optimization horizon) (Metz et al., 2021). Evolutionary methods can help mitigate this because they do not rely directly on gradients, and thus are less susceptible to gradient sensitivity issues. However, the complexity of the problem increases with the length of the optimization horizon. Short horizon bias (Wu et al., 2018) constitutes another problem, where meta-optimization over a specified time horizon introduces a bias that prevents it to be applicable to a longer time horizon.
Meta-generalization refers to a learned optimizer’s ability to perform well on novel tasks it was not specifically trained for. One appraoch to achieving this is to train the optimizer on the widest possible range of tasks. Previous research focused on developing a single, adaptable algorithm by training it on a very broad domain. For example, Versatile Learned Optimizers (VeLO, Metz et al., 2022) was trained on hundreds of different problems, ranging from simple linear regression to reinforcement learning. However, this “one-size-fits-all” approach has significant drawbacks. The computational cost of training such an optimizer is immense, requiring 4 00040004\,0004 000 TPU-months in the case of VeLO. Furthermore, a single optimizer struggles to effectively adapt to the vast array of loss surfaces encountered across such a diverse set of tasks. This makes it challenging to achieve consistently high performance across all tasks.
Instead of aiming for broad applicability, we focus on the increasingly relevant domain of fine-tuning pretrained models. This specialization allows us to develop learned optimizers in a more targeted way. By narrowing the scope of the meta-training domain, the learned optimizer can specialize and excel in the specific types of tasks it is designed for. Fine-tuning also simplifies the optimization problem by leveraging existing knowledge encoded in pretrained checkpoints (Wei et al., 2021) and is typically done for a relatively few iteration, since only a small alignment with the current task is needed to achieve a good results.
To achieve this, we introduce L3RS (Learned Layer-wise Learning Rate Scheduler, pronounced “lers”), a novel learned optimizer designed specifically to leverage the performance of a set of predefined base optimizers within a narrower domain. L3RS uses a variety of performance features, such as adaptive exponential moving averages (EMA) across different scales. EMAs, or model averaging, have demonstrated improved generalization in real-world applications (Tarvainen & Valpola, 2017; Izmailov et al., 2018). Based on these features, L3RS produces layer-wise updates as a linear combination of the predefined base optimizers, along with a corresponding learning rate for each layer.
Our contributions are as follows.
  • We propose a novel learning-to-learn method that leverages the strengths of multiple base optimization methods. This optimizer is simple and intuitive, having only a fraction of parameters compared to other black-box learned optimizers.
  • We demonstrate that meta-training a learned optimizer on a narrow domain results in an optimizer that can outperform existing methods, with more that 50% speedup compared to another learned optimizer baseline, and more than 2×2$\times$2 × speedup compared to best performing traditional optimizers.
  • We evaluate the meta-generalization of the proposed optimizer across several components: generalization to longer training horizon, different pretraining and model initializations, and different evaluation dataset.
2 Preliminaries
[…]
Figure 1: Left: Inner loop evaluation. Given a task $T:=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$, the learned optimizer $f_{\psi}$ uses optimization statistics $\Phi$ to update model parameters starting from $\theta_{0}$ for each of $\{D^{t}_{K}\}$ batches. After $K$ update steps, the final model parameters are evaluated on the evaluation set $D^{e}$ using meta-loss $L_{\psi}$. Right: Outer loop NES meta-training iteration. Given a task $T$ and meta-parameters $\psi_{T}$, Gaussian noise is added to $\psi_{t}$ to produce a number of candidates equal to the population size $c$, $(\psi_{t,0},...,\psi_{t,c})$. An inner loop evaluation is performed on the given task for all candidates. The fitness of each candidate is then used to perform an NES update step, resulting in the next learned optimizer parameters, $\psi_{t+1}$.
Consider a loss function $L(\theta,X)$, parameterized by $\theta$ for a given data $X$. Learned optimizers $f_{\psi}$, parameterized by $\psi$, provide a direction to update the model’s weights $\theta$ as
$\theta_{n+1}=\theta_{n}+f_{\psi}(\Phi;\theta_{n},X),$ (1)
where $\Phi$ is a collection of features that describes the current (or past) optimization statistics, such as the gradient or the loss value. For example, the simplest learned optimizer might simply optimize the learning rate $\lambda$ of SGD, as $f_{\psi}(\Phi;\theta_{n},X)=\lambda\nabla L(\theta_{n},X)$, in which case $\Phi:=\{\nabla L\}$ and $\psi:=\{\lambda\}$. In case of Adam, the optimization parameters would be $\psi:=\{\lambda,\beta_{1},\beta_{2}\}$.
In Figure 1 we show both the inner and the outer loop of a typical meta-training process. We train $\psi$ on a distribution of tasks from a given dataset. A task is a set $T:=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$, where $\theta_{0}$ represents initial model weights (or a distribution of initial model weights), $\{D^{t}_{K}\}$ is a sequence of $K$ training batches, and $D^{e}$ represents an evaluation batch. A batch typically consists of data $X$ and, for classification problems, may also include labels. We assume that $\theta_{0}$ is given as input, and we have no control over its generation process (unlike initialization-based meta-learning approaches, such as MAML, Finn et al., 2017).
Learned optimizer parameters $\psi$ are evaluated by performing $K$ update steps (1) on every batch from $D^{t}_{K}$, to get $\theta_{K}$ and calculating the loss using the evaluation batch $L_{\psi}(\theta_{K},D^{e})$.
We use Natural Evolution Strategies (NES, Salimans et al., 2017) to meta-train the learned optimizers. We define a task meta-loss $L_{\psi}^{M}(T)=L_{\psi}(\theta_{K},D^{e})$, the loss on the evaluation batch after K training steps of a task $T=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$. We then define a fitness function $F(\psi,(T_{1},...,T_{b}))=-\frac{1}{b}\sum^{b}_{i=1}L_{\psi}^{M}(T_{i})$, using a batch of $b$ tasks for each generation. NES is used to maximize the fitness function during meta-training.
After the learned optimizer is trained, it can be evaluated to a new task distribution $\widetilde{T}:=\{\widetilde{\theta}_{0},\{\widetilde{D}^{t}_{K}\},\widetilde{D}^{e}\}$.
3 Motivation and Related Work
We can separate the optimizers into two main categories. First category, so called black-box optimizers (Li & Malik, 2016; Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017; Sandler et al., 2021; Metz et al., 2020a, b, 2022), learns an update function $f_{\psi}$ from scratch using a custom inner optimization loop. These methods often have a large number of parameters, making them prone to overfitting and stability issues (Harrison et al., 2022). While recent work has explored using Transformers as a meta-learner architecture (Chen et al., 2022; Moudgil et al., 2023; Jain et al., 2024), they have not yet shown a significant advantage over traditional optimizers or other learned optimizer architectures.
A notable example of black-box optimizer is VeLO, which uses per-tensor HyperNetworks (Ha et al., 2016) represented by 512-wide LSTMs to generate parameters of per-parameter MLPs. To compute the parameter updates, it passes the per-parameter features through the generated MLP network consisting of 2-hidden layer, 4-hidden MLP. The input to the HyperNetwork and MLP vary slightly, but generally represent the training dynamic of the optimization. VeLO is designed to be general and has a lot of parameters that need to be learned, which makes meta-training very expensive. We envision an ideal optimizer would be lean and ideally fast to train on a narrow task domain.
Other techniques utilize algorithm discovery as a program search (Bello et al., 2017; Wang et al., 2022; Zheng et al., 2022; Chen et al., 2023), where the symbolic optimizers are found using a tree search of predefined operations (such as gradient and momentum).
The second category of meta-optimizers learns a higher-level meta-component on top of existing hand-designed optimizers. Such approaches include learning to adapt the hyperparameters of an underlying optimizer Daniel et al. (2016), or learn a schedule for the learning rate (Xu et al., 2017, 2019). In addition, these methods can include adapting the optimizer’s parameters directly (Shaban et al., 2019). Hypergradient methods (Maclaurin et al., 2015; Baydin et al., 2017; Grazzi et al., 2020; Moskovitz et al., 2019) perform a hyperparameter search on-the-fly during optimization.
Our proposed approach strikes a balance between flexibility of black-box minimizer and the stability of learning hyperparameters of known optimizers. We leverage the benefits of hand-designed optimizers while incorporating the adaptability of learned components using a simple small neural network for learning the per-layer learning rate. Unlike existing methods that learn a fixed set of hyperparameters for a single existing optimizer, L3RS learns from a set of base-optimizers. Moreover, during meta-training we also search over the hyperparameters of the base-optimizer.
Almeida et al. (2021) is probably the closest algorithm to our setting. It uses an LSTM controller to adjust the hyperparameters of a hand-designed inner-optimizer, which uses a collection of hand-designed parts from different known optimizers. A major difference with our method is that L3RS operates per-layer and that our algorithm wraps existing optimizers out-of-the box, without the need to manually add features.
Prémont-Schwarz et al. (2022) also propose an in-between approach, where the learned optimizer falls back to the base-optimizer. However, their goal was to ensure convergence, not necessarily to improve the performance.
The L3RS architecture was also partially motivated by Learning Rate Grafting (Agarwal et al., 2022), which demonstrates the utility of isolating the direction of an optimizer from its magnitude.
Similar to our paper Landro et al. (2021) also propose an algorithm that combines Adam and SGD as a linear combination of their update direction. However, their method differs significantly in two key aspects: (1) they do not provide a mechanism for learning the mixing coefficient, instead treating it as fixed hyperparameters, and (2) their combination is applied globally across all layers, while our proposed method, L3RS, employs a layer-wise adaptive strategy.
4 L3RS: Learned Layer-wise Learning Rate Scheduler
Figure 2: Left: L3RS applied to a single layer of the target model. The MLP receives the norm of each direction provided, time features, target layer embedding, and EMA features of the global loss, layer gradient norm and layer parameter norm. The MLP outputs the weighting for each direction ($\mu_{p}$) as well as the final update norm ($\lambda$). Right: L3RS is applied at every layer of the network independently with the same MLP weights but different layer embeddings.
4.1 Model architecture
Given a set of update vectors $d_{p}$ from $P$ input base optimizers (e.g. SGD or Adam) L3RS computes a model parameter update step per layer $l$ using:
${\Delta\theta^{(l)}=\lambda^{(l)}\sum_{p=1}^{P}\mu_{p}^{(l)}\hat{d}_{p}^{(l)},}$ (2)
where $\lambda^{(l)}$ is a layer-specific update $l_{2}$-norm, $\mu_{p}^{(l)}$ are normalized mixing coefficients for the base optimizers’ directions $\sum_{p=1}^{P}\mu_{p}^{(l)}=1$, and $\hat{d}_{p}^{(l)}$ are normalized optimizers’ direction $\hat{d}_{p}^{(l)}=\frac{d_{p}^{(l)}}{\left\lVert d_{p}^{(l)}\right\rVert_{2}}$.
Figure 2 shows the main input and output components of L3RS as well is provide a flow-chart on how the weights of the underlying model $\theta$ are updated.
In order to learn $\lambda^{(l)}$ and $\mu_{p}^{(l)}$, we provide L3RS with a feature vector, comprising the adaptive EMA, time, embedding features as well as the magnitude of the update directions defined below. This input is then processed by two fully connected layers with sizes $32$ and $16$, using ReLU activation. The model outputs logits $z^{(l)}\in\mathbb{R}^{P+1}$, from which we reconstruct $\mu_{p}^{(l)}=\frac{\exp(z_{p}^{(l)})}{\sum_{j=1}^{P}\exp(z_{j}^{(l)})}$ and $\lambda^{(l)}=\exp(z_{P+1}^{(l)})$. The hyperparameters of the input optimizer (e.g. Adam’s $\beta_{1}$ and $\beta_{2})$) are also jointly meta-learned along with the L3RS’s model parameters. This enables co-adaptation of the layer-wise scaling and the underlying optimization algorithm.
Adaptive Exponential Moving Averages (EMA).
In order to capture both short-term and log-term performance trends, for each layer $l$, we maintain four EMAs with different smoothing factors $\gamma\in\{0,0.9,0.99\}$. EMAs are updated recursively as $a^{(l)}_{i,k+1}:=\gamma_{i}a^{(l)}_{i,k}+(1-\gamma_{i})\xi^{(l)}$, where $a^{(l)}_{i,k}$ denotes the $i$th EMA for layer $l$ at time step $k$, and $\xi^{(l)}$ represents a layer-specific statistic of interest. Specifically, for a given layer we track the following:
  • Log $l_{2}$-norm of weights. $\log(\left\lVert w^{(l)}\right\rVert_{2})$, where $w^{(l)}$ are the weights for layer $l$..
  • Log $l_{2}$-norm of gradients. $\log(\left\lVert g^{(l)}\right\rVert_{2})$, where $g^{(l)}$ are the gradients for layer $l$.
  • Loss. The loss of the overall model (shared across all layers).
To ensure unbiased estimates, especially during the initial time steps, we apply a bias correction to the EMAs before incorporating them as input features: $\tilde{a}^{(l)}_{k}=\frac{a^{(l)}_{k}}{1-\gamma^{k}}.$
Time features.
Inspired by VeLO (Metz et al., 2022), we propose incorporating the following time features for the explicit modeling of temporal dynamics of the optimization progress. These features enable the optimizer to adapt its strategy based on the stage of the training process. These features are generated using $k$, the current training step, and $K$, the number of total training steps. Similar to some learning rate schedulers, $K$ must be given to the optimizer at the beginning of training so that these time features can be generated.
  • Relative time features. $\tanh\left(10\left(\frac{k}{K}-\alpha_{i}\right)\right)$, where $\alpha_{i}$ are $11$ linearly scaled reference points between $0.0$ and $1.0$, inclusive. These features capture the relative progress through a sequence, with the hyperbolic tangent function providing a smooth, bounded representation. Figure 7 in the Appendix shows the values of these features throughout training.
  • Absolute time features. $\tanh\left(\log\left(K\beta_{i}\right)\right)$, where $\beta_{j}$ are $4$ log-scaled scaling factors between $0.0001$ and $0.1$. These features provide a logarithmic encoding of the total duration, which can be useful for appropriate learning rate scaling. For longer time horizons these features can be extended such that $\max(\beta)>\max(K)$. Figure 7 in the Appendix shows the values of these features for various total training lengths $K$.
Embedding features.
A 16-dimensional embedding vector $\psi_{emb}$ is meta-learned for each layer, allowing the optimizer to learn specialized per-layer dynamics.
Base-optimizer direction magnitude.
For each base optimizer $d_{p}$, the log $l_{2}$-norm of the direction is provided: $\log(\left\lVert d_{p}\right\rVert_{2})$.
4.2 Comparison with VeLO
Optimizer State Memory Update Function Size
SGD $\times0$ 61.4 KiB
Adam 2.0×2.0$\times$2.0 × 240.0 KiB
L3RS 2.0×2.0$\times$2.0 × 2.9 MiB
VeLO 4.0×4.0$\times$4.0 × 418.78 MiB
Table 1: Compute and memory costs for various optimizers. State memory is the ratio of the optimizer state size to the model parameter size. Update Function Size is the size of the compiled code to calculate the weight update given the gradients. Target model is ResNet-34 with 1000 classes.
VeLO was designed with an efficient hypernetwork-style architecture. The majority of the computation cost is reduced to per-layer LSTM networks, which then generate cheaper per-parameter networks which output the update step direction. L3RS is able to leverage smaller MLP networks for the per-layer operations, and can rely on base optimizers rather than generating the update step direction manually. These design choices result in a learned optimizer which has approximately 620×620$\times$620 × fewer parameters.
5 Experiments
To evaluate the performance of our proposed L3RS optimizer, we designed a series of image classification experiments using the ResNet-34 model.
Model Choices.
While our proposed L3RS can wrap any number of base-optimizers, we found that even a simple version that combines together only Adam and SGD (without momentum) directions performs well. For the remainder of this paper, when we refer to L3RS, we specifically mean this Adam and SGD combination unless we specify otherwise.
During inference, the memory footprint of this L3RS optimizer state is 2×2$\times$2 × the model parameters (all from Adam). VeLO uses a memory state of 4×4$\times$4 × the model parameters (plus additional memory from $3$ AdaFactor style accumulators). Table 1 shows a comparison of the memory state requirements for various optimizers, as well as the compute cost for each.
We derive updates separately for every convolutional, dense or batch normalization layer of the model, including separate updates for kernels and biases. For ResNet-34, this results in $L=111$ components for which we compute learning rate $\lambda^{(l)}$ and mixing coefficients $\mu_{p}^{(l)}$, for $l=1,\dots,L$. With $P=2$ optimizers, L3RS outputs $333$ parameters per iteration.
For NES training, we use a smooth exponential decay for both meta-learning rate $\alpha$ and the Gaussian noise standard deviation $\sigma$ by 0.5 every 500 generations. We use the same antithetic sampling and fitness transformations as in Salimans et al., 2017. We use a population size of $32$, meta-batch size of $4$ and train for $2000$ generations for all our experiments (except the ablation study). We parallelize the evaluation of each candidate in the population, using a total of $32$ A100 GPUs for four days.
Task Distribution and Meta-Training.
We trained the learned optimizer using ImageNet dataset in the following way. We partition the ImageNet dataset into three distinct subsets: pretraining, meta-train (ImageNet25), and meta-test (ImageNet25Eval). The first $500$ classes are used to pretrain a model using a conventional off-the-shelf algorithm. The pretrained model serves as an initialization to our method. The remaining $500$ classes are randomly divided into meta-train and meta-test sets. Details on checkpoint pretraining are provided in Appendix A.
A task within our meta-learning framework is constructed by randomly sampling $25$ classes from either the ImageNet25 or ImageNet25Eval set. The selected classes are used to generate $K$ training batches with $64$ samples each and $1$ evaluation batch with $256$ samples, using the train and validation splits respectively. During meta-training, the number of training batches $K$ is uniformly sampled from a range of $10$ to $500$, simulating diverse fine-tuning scenarios.
We meta-train the L3RS optimizer and the VeLO optimizer on the ImageNet25 meta-train task distribution using NES as described above.
Meta-Evaluation and Baselines.
To evaluate the performance, we check for in-distribution and out-of-distribution generalization. For in-distribution evaluation we use the same initialization as the meta-training ($\widetilde{\theta}_{0}=\theta_{0}$) and use ImageNet25Eval tasks for meta-test. For out-of-distribution evaluation we use the Places (López-Cifuentes et al., 2020) dataset. We create a separate Places initialization by pretraining ResNet-34 on the first $150$ classes of Places. We use the remaining classes as a meta-test set (Places25Eval). Importantly, this dataset is entirely novel and unseen during the meta-training, which allows us to evaluate out-of-distribution performance of the optimizers.
A $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Refer to caption B $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: Places25Eval
Refer to caption C $\widetilde{\theta}_{0}$: Places,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Refer to caption D $\widetilde{\theta}_{0}$: Places,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: Places25Eval
Refer to caption E $\widetilde{\theta}_{0}$: RandomInit,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Refer to caption F
Refer to caption Refer to caption
Figure 3: Meta-evaluation of the L3RS, meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks. Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right). Each marker represents model evaluation at that number of steps. Solid lines indicate the number of steps for in-distribution evaluation, while dashed lines indicate generalization to more steps than meta-training.
A. In-domain Generalization. Both initialization and evaluation are on ImageNet.
B. Out-of-Domain Initialization. Initialized on ImageNet, evaluated on Places25Eval.
C. Out-of-Domain Evaluation. Initialized on Places, evaluated on ImageNet25Eval.
D. Out-of-Domain Init & Eval. Both initialization and evaluation are on Places dataset.
E. Random Initialization. Random initialization, evaluated on ImageNet25Eval dataset.
F. Speedup of L3RS. For in-domain generalization, this shows how much faster L3RS achieves a given accuracy compared to the baselines.
We compare L3RS against the following baselines:
  • VeLO (ft). An instance of VeLO meta-trained using the same method as L3RS.
  • VeLO (og). The original pretrained VeLO model (Metz et al., 2022), representing a general-purpose learned optimizer.
  • VeLO (og, Head). The original pretrained VeLO model applied to only the final model layer (freezing the rest of the model).
  • Adam (cosine). Adam optimizer with a cosine learning rate schedule.
  • Adam (cosine, Head). Adam optimizer with a cosine learning rate schedule applied to only the final model layer (freezing the rest of the model).
  • Adam (const). Adam optimizer with constant learning rate.
  • Adam (const, Head). Adam optimizer with constant learning rate appled to only the final model layer (freezing the rest of the model).
For optimizers which parameters depend on the number of training steps $K$, such as Adam (cosine), VeLO (ft) and L3RS, we perform separate evaluations across various $K$ values to assess their performance both within and beyond the meta-training regime.
For all meta-evaluations we average the results of 100 tasks sampled from the specified task distribution. The average and standard deviation of eval accuracy across the sampled tasks is reported.
Results.
Figure 3 shows the main result. Our goal is to assess the performance of the learned optimizer under various conditions, for both in and out of distribution scenarios. (A) Represents in-distribution evaluation, where the initialization, evaluation dataset, and the number of steps match the meta-training regime. We also explore out-of-distribution performance by extending the number of evaluation steps beyond the meta-training range (indicated by the dashed line in all plots), changing the evaluation dataset to Places25Eval (B), modifying the initialization to utilize a model pretrained on Places (C), or using both the Places checkpoint and Places25Eval dataset (D).
Figure 4: Learned mixing coefficients $\mu_{p}^{(l)}$ and $\lambda^{(l)}$ for each iteration of L3RS for ResNet-34 model. Top: average across all the layers. Bottom: individual for every learned component.
First, comparing VeLO (ft) to VeLO (og), we observe that fine-tuning VeLO to a specific dataset and number of steps leads to a significant performance improvement compared to the general-purpose VeLO (og). However, VeLO (ft) performance deteriorates sharply when evaluated outside its meta-training regime. We hypothesize this is due to VeLO complexity and parameter count, making it more sensitive to deviations from its training distribution, especially number of steps.
[…]
Interestingly, changing either the initialization or evaluation dataset to Places has a minimal effect on the relative perfoamnce of the optimizers, even though Places was not included in the meta-training data. This highlights the robust generalization capabilities of both L3RS and the fine-tuned VeLO.
We also apply the methods to a randomly initialized Resnet-34 model (E). RandomInit uses a standard initialization method (Klambauer et al., 2017). In this case, L3RS does not outperform the baseline methods. We believe this is because, while L3RS generalizes well to different checkpoints (C), the randomly initialized model may be too different and far out of distribution for L3RS to perform optimally.
In Figure 3 (F) we show the speedup, in training steps, L3RS achieves over baseline methods to reach a given accuracy. L3RS robustly is able to demonstrate 50% speedup over VeLO (ft) and $100\%$ speedup over the best hand-designed optimizer.
Finally, Figure 4 shows the learned parameters of L3RS during a $100$-step evaluation run. Interestingly, the optimizer initially heavily favors Adam direction, switching towards SGD as training progresses (with a curious sudden switch for some layers around iteration $10$). Moreover, some layers consistently rely on SGD or Adam direction throughout the entire run. While the precise interpretation of these parameter dynamics is challenging, the parameters of L3RS are considerably more interpretable than those of most black-box learned optimizers.
Base-optimizers Global No Embedding Per-layer MLP Embedding
SGD only $-2.26pm0.20$ $64.48\pm 5.05$ $64.71\pm 5.00$ $68.18\pm 4.29$
Adam only $-5.36pm0.20$ $61.44\pm 5.14$ $67.14\pm 4.57$ $68.52\pm 4.25$
SGD, Adam $56.71\pm 4.93$ $65.70\pm 4.91$ $68.52\pm 4.40$ $\mathbf{68.93\pm 4.58}$
Table 2: Average and standard deviation evaluation accuracy given different optimizers and per-layer strategies.
EMA’s Smoothing Factors Accuracy
$\gamma\in\{0.99,0.9,0.0\}$ $68.93\pm 4.34$
$\gamma\in\{0.9,0.0\}$ $67.43\pm 4.64$
$\gamma\in\{0.0\}$ $67.78\pm 4.47$
$\gamma\in\varnothing$ $66.52\pm 4.84$
Table 3: Average and standard deviation evaluation accuracy given adaptive EMA features.
6 Ablations and Variants
We perform a methodical ablation and variant study to justify the design choices of the L3RS architecture. We meta-train all models in this section using the ImageNet25 task distribution but set number of steps, $K$, to $10$. We additionally meta-train for $500$ generations of NES and change the meta-learning rate decay rate steps from $500$ to $100$.
Ablation of Base Optimizers.
We first show the results of using either SGD or Adam alone as the wrapped optimizer rather than both together.
Embedding Variants.
Without providing layer embeddings to L3RS, the optimizer would have more difficulty identifying which layer it was operating on needing to rely heavily on the adaptive EMA features. An alternative to the layer embeddings is to meta-learn a separate MLP model for each layer of the target model. This strategy greatly increases the total parameter count of the optimizer. We investigate both the per-layer MLP style, as well as the single MLP without layer embeddings. We additionally report results for a Global method, which applies a single learning rate, $\lambda$, and direction weights, $\mu_{p}$, for all layers in the model. See Table 2 for results.
Ablation of Adaptive EMA Input Features.
We additionally train models with different adaptive EMA feature smoothing factors. While L3RS uses $\gamma\in\{0.99,0.9,0.0\}$, we also test $\gamma\in\{0.9,0.0\}$, $\gamma\in\{0.0\}$ (which corresponds to the raw features without averaging), and $\gamma\in\varnothing$ (which corresponds to not receiving the adaptive features at all). See Table 3 for results.
Base Optimizer Set Accuracy
SGD, Adam $\mathbf{68.93\pm 4.58}$
Adam, Lion, LAMB $68.61\pm 4.50$
Adamax, SGD, LAMB $66.95\pm 4.80$
SGD, Adam, Adamax, Lion, LAMB, WeightDecay $68.41\pm 4.65$
Table 4: Average eval accuracy and standard deviation of L3RS variants, each using a different set of base-optimizers.
New Directions.
Our main results are based on L3RS which uses only SGD and Adam as given directions. There is a lot of room for exploration in which optimizers and combinations will result in the most powerful L3RS variant for a given task. As an example of the flexibility of the architecture we explore a few combinations here. We leverage the following optimizers: Lion (Chen et al., 2023), LAMB (You et al., 2020), Adamax (Kingma & Ba, 2014), as well as a WeightDecay direction which simply provides the negative direction of the current parameters. We use the same training/evaluation set up as the rest of the ablations and variants. We report the results in Table 4.
7 Conclusions
Learned optimization is a powerful paradigm which has the potential to outperform existing methods. Our results suggest that narrowing the domain that learned optimizers are meta-trained on can provide a reduced meta-training cost, and allow learning domain specific exploitations that can increase their performance, especially on fine-tuning tasks. We provide results repurposing VeLO, an architecture designed for general learned optimization, as well as propose a new architecture L3RS, designed to take advantage of the narrow domain. We also demonstrate the robustness of this method on out of distribution tasks. As fine-tuning tasks continue to become more relevant, especially for LLMs, this paradigm and architecture provides a promising direction for future research. Improving this technique further can result in faster model training, better performance, and robust data generalization.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims in the abstract are: exploring training learned optimizers for a narrow task domain, proposing a new architecture, the evaluation methods, and a brief description of baselines. The experiments justifying these claims are explained in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The scope of the domain in the paper is purposefully limited, the idea being to train on an extremely narrow task distribution. We do show experiments evaluating outside of this distribution, some of which perform well and others which do not. We propose explanations for why the method generalizes to some distributions but not others. We also acknowledge that while we only used Adam and SGD as directions provided to L3RS for the main training runs, we give examples and some results of using others.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not include any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: There is all information necessary to replicate the experiments in the paper. Code is not provided but the datasets used are public. The architecture proposed is sufficiently explained to reproduce without much difficulty. We intend to release the code after the paper’s acceptance.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: There is all infomation necessary to replicate the experiments in the paper. Code is not provided but the datasets used are public. The architecture proposed is sufficiently explained to reproduce without much difficulty.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We’ve ensured to specify every hyperparameter for meta-training and task generation used to facilitate reproducibility and understanding.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: When applicable error bars (for graphs) and standard deviations (for tables) are reported.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe the computational power required to perform out meta-training. The evaluation computation requirements are not specified, but is relatively cheap and understandable (typical Resnet-34 training requirements).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper uses only well established, public datasets.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This method establishes a new learned optimizer architecture and type of training domain, neither of which have any obvious societal implications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not release any models and only uses public datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Datasets and baseline methods used are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This work does not introduce any assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Checkpoint Pretraining
We create the pretrain checkpoints for both ImageNet and Places using the same hyperparameters. We train the models using batch size $128$ and train for $100,000$ steps. Augmentations/preprocessing used include random cropping to size 224, random mirror, resize, and normalization (based on Imagenet train statistics). For the ImageNet checkpoint trained on the first 500 classes, we reach an evaluation loss of $0.9387$ and an evaluation accuracy of $75.68$. For the Places checkpoint trained on the first 150 classes, we reach an evaluation loss of $1.312$ and an evaluation accuracy of $61.72$.
Appendix B Meta-Training
We provide the meta-training curves, both fitness and accuracy, for the two meta-trained learned optimizers. Figure 5 shows the curves for L3RS and VeLO throughout training. VeLO training is very unstable in this setup, though does converge by the end of training. L3RS reaches a high fitness very rapidly then slowly improves throughout the rest of the training, likely due to the smaller number of parameters in the learned optimizer.
Figures 7 and 7 show the example of the relative and absolute time features used by L3RS.
Appendix C Evaluation Tables
We provide the full evaluation tables from the main evaluation experiments Figure 5 A-E. Table 5 provides the results from the in-domain evaluation A. Table 6 provides the results for all head-only fine-tuning for evaluation A. Table 7 provides the results from the dataset-generalization evaluation B. Table 8 provides the results from the checkpoint-generalization evaluation C. Table 9 provides the results from the initialization and dataset-generalization evaluation D. Table 10 provides the results from the RandomInit evaluation E.
Figure 5: Meta-training curves for L3RS (Black), and VeLO (Blue). Left: Average population fitness. Right: Average population evaluation accuracy.
Figure 6: Relative time features as a function of training progress (k/K). Each of the lines represents a time input feature to the model, generated using $k$, $K$, and $\alpha_{i}$.
Figure 7: Absolute time features as a function of max training step (K). Each of the lines represents a time input feature to the model, generated using $K$ and $\beta_{j}$.
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
L3RS $68.52\pm 4.46$ $76.78\pm 3.83$ $80.22\pm 3.45$ $83.54\pm 2.99$ $86.57\pm 2.77$ $88.47\pm 2.64$ $88.79\pm 2.57$ $89.44\pm 2.52$
VeLO (ft) $62.11\pm 5.51$ $73.23\pm 4.29$ $77.41\pm 3.44$ $81.54\pm 3.46$ $85.16\pm 2.93$ $86.64\pm 3.1$ $73.64\pm 26.26$ $4.0\pm 0.17$
VeLO (og) $50.63\pm 4.9$ $47.48\pm 5.2$ $53.7\pm 4.6$ $60.79\pm 4.24$ $71.02\pm 4.08$ $78.09\pm 3.62$ $82.45\pm 3.19$ $83.39\pm 3.04$
Adam (cosine), 1e-3 $55.86\pm 4.92$ $66.38\pm 4.7$ $71.03\pm 4.11$ $76.59\pm 4.19$ $82.53\pm 3.46$ $85.34\pm 3.17$ $86.84\pm 2.7$ $87.25\pm 2.88$
Adam (cosine), 1e-4 $10.69\pm 2.52$ $28.22\pm 4.03$ $52.11\pm 5.04$ $70.29\pm 4.6$ $81.22\pm 3.43$ $85.3\pm 3.22$ $87.52\pm 2.96$ $88.88\pm 2.59$
Adam (cosine), 1e-5 $4.32\pm 1.45$ $5.09\pm 1.62$ $6.47\pm 1.81$ $10.66\pm 2.29$ $29.4\pm 4.0$ $54.18\pm 5.05$ $72.22\pm 4.67$ $82.36\pm 3.52$
Adam (const), 1e-3 $60.12\pm 4.62$ $62.65\pm 4.37$ $65.82\pm 4.24$ $70.85\pm 4.23$ $75.92\pm 3.94$ $79.11\pm 3.7$ $80.87\pm 3.58$ $81.92\pm 3.61$
Adam (const), 1e-4 $22.09\pm 3.54$ $51.72\pm 4.98$ $69.72\pm 4.59$ $78.58\pm 3.95$ $84.02\pm 3.07$ $86.65\pm 2.93$ $87.72\pm 2.76$ $87.95\pm 2.78$
Adam (const), 1e-5 $4.83\pm 1.56$ $6.47\pm 1.77$ $10.7\pm 2.29$ $22.66\pm 3.62$ $53.91\pm 4.99$ $72.15\pm 4.6$ $80.34\pm 3.62$ $85.62\pm 3.16$
Table 5: Average and standard deviation evaluation accuracy for main experiment (A).
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
VeLO (og) $42.53\pm 4.73$ $67.93\pm 4.4$ $73.68\pm 3.65$ $76.35\pm 3.58$ $79.12\pm 3.45$ $80.71\pm 3.2$ $81.36\pm 3.11$ $81.6\pm 3.23$
Adam (cosine), 1e-3 $7.98\pm 2.16$ $16.93\pm 3.14$ $34.58\pm 4.39$ $54.65\pm 5.05$ $69.84\pm 4.55$ $74.95\pm 4.0$ $78.0\pm 3.68$ $80.52\pm 3.41$
Adam (cosine), 1e-4 $4.18\pm 1.47$ $4.53\pm 1.59$ $5.32\pm 1.7$ $7.52\pm 2.06$ $16.86\pm 3.24$ $34.34\pm 4.53$ $54.8\pm 4.99$ $70.14\pm 4.72$
Adam (cosine), 1e-5 $3.95\pm 1.38$ $3.98\pm 1.38$ $4.05\pm 1.39$ $4.21\pm 1.45$ $4.55\pm 1.54$ $5.24\pm 1.55$ $7.32\pm 1.96$ $16.74\pm 3.15$
Adam (const), 1e-3 $13.31\pm 2.71$ $33.5\pm 4.31$ $53.59\pm 4.81$ $65.91\pm 4.67$ $74.17\pm 4.07$ $77.34\pm 3.67$ $79.53\pm 3.51$ $81.17\pm 3.45$
Adam (const), 1e-4 $4.44\pm 1.51$ $5.22\pm 1.59$ $7.4\pm 1.97$ $13.27\pm 3.03$ $34.17\pm 4.59$ $54.5\pm 4.94$ $67.28\pm 4.72$ $75.17\pm 3.93$
Adam (const), 1e-5 $3.98\pm 1.38$ $4.01\pm 1.41$ $4.21\pm 1.43$ $4.43\pm 1.52$ $5.22\pm 1.57$ $7.32\pm 2.0$ $13.13\pm 3.05$ $34.35\pm 4.57$
Table 6: Average and standard deviation evaluation accuracy for all Head-only fine-tuning from main experiment (A).
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
L3RS $59.2\pm 3.87$ $69.16\pm 3.32$ $73.64\pm 3.23$ $76.75\pm 3.32$ $80.67\pm 3.07$ $82.74\pm 2.92$ $84.32\pm 2.82$ $85.51\pm 2.81$
VeLO (ft) $50.13\pm 7.32$ $64.78\pm 3.46$ $70.46\pm 3.29$ $74.1\pm 7.83$ $79.12\pm 3.21$ $76.54\pm 12.9$ $56.13\pm 30.47$ $4.02\pm 0.36$
VeLO (og) $43.13\pm 3.98$ $44.58\pm 4.18$ $51.26\pm 4.18$ $59.02\pm 4.0$ $68.38\pm 3.83$ $75.14\pm 3.47$ $79.97\pm 3.49$ $83.07\pm 3.13$
Adam (cosine), 1e-3 $48.85\pm 3.79$ $60.39\pm 3.34$ $66.11\pm 3.59$ $71.81\pm 3.64$ $77.71\pm 3.44$ $80.73\pm 3.17$ $82.99\pm 2.96$ $84.97\pm 2.76$
Adam (cosine), 1e-4 $9.87\pm 2.24$ $23.3\pm 2.97$ $44.16\pm 3.99$ $63.11\pm 3.72$ $74.9\pm 3.65$ $79.71\pm 3.0$ $82.68\pm 3.04$ $85.17\pm 2.98$
Adam (cosine), 1e-5 $4.38\pm 1.53$ $4.98\pm 1.55$ $6.07\pm 1.68$ $9.06\pm 2.06$ $22.83\pm 3.06$ $45.84\pm 3.56$ $65.3\pm 3.72$ $76.39\pm 3.54$
Adam (const), 1e-3 $52.54\pm 4.3$ $56.19\pm 3.93$ $61.15\pm 3.92$ $65.87\pm 3.94$ $71.71\pm 3.52$ $74.83\pm 3.42$ $77.42\pm 3.57$ $80.51\pm 3.28$
Adam (const), 1e-4 $17.32\pm 2.81$ $43.3\pm 3.74$ $62.03\pm 3.48$ $71.65\pm 3.48$ $78.01\pm 3.28$ $80.57\pm 2.85$ $82.72\pm 3.01$ $84.66\pm 2.97$
Adam (const), 1e-5 $4.73\pm 1.54$ $6.08\pm 1.66$ $9.05\pm 2.15$ $17.72\pm 2.64$ $45.84\pm 3.79$ $64.93\pm 3.85$ $74.3\pm 3.65$ $80.72\pm 3.3$
Table 7: Average and standard deviation evaluation accuracy for main experiment (B).
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
L3RS $65.65\pm 4.61$ $73.13\pm 3.99$ $77.0\pm 4.02$ $80.33\pm 3.59$ $83.79\pm 3.06$ $85.93\pm 2.97$ $86.91\pm 2.78$ $87.04\pm 2.68$
VeLO (ft) $57.85\pm 7.39$ $69.6\pm 4.0$ $73.51\pm 4.09$ $77.96\pm 3.8$ $82.18\pm 3.22$ $84.46\pm 2.95$ $82.67\pm 11.37$ $4.04\pm 0.27$
VeLO (og) $52.35\pm 4.67$ $49.56\pm 4.78$ $54.52\pm 4.72$ $61.44\pm 4.58$ $71.37\pm 4.1$ $77.87\pm 3.67$ $81.7\pm 3.15$ $82.82\pm 3.08$
Adam (cosine), 1e-3 $55.36\pm 4.8$ $62.93\pm 4.44$ $67.45\pm 4.35$ $72.54\pm 4.18$ $79.14\pm 3.41$ $82.61\pm 3.09$ $84.43\pm 3.19$ $85.36\pm 3.17$
Adam (cosine), 1e-4 $13.27\pm 2.7$ $33.25\pm 4.39$ $54.11\pm 4.84$ $68.21\pm 4.46$ $78.83\pm 3.61$ $83.49\pm 3.16$ $85.96\pm 2.98$ $87.31\pm 2.66$
Adam (cosine), 1e-5 $4.37\pm 1.52$ $5.26\pm 1.72$ $7.14\pm 2.02$ $12.29\pm 2.7$ $33.43\pm 4.16$ $56.37\pm 4.73$ $70.46\pm 4.56$ $80.28\pm 3.51$
Adam (const), 1e-3 $56.2\pm 4.9$ $58.38\pm 4.31$ $61.7\pm 4.4$ $66.41\pm 4.45$ $72.24\pm 4.05$ $75.95\pm 3.68$ $78.45\pm 3.38$ $80.04\pm 3.6$
Adam (const), 1e-4 $25.4\pm 3.79$ $53.72\pm 4.75$ $67.04\pm 4.56$ $75.64\pm 4.01$ $81.56\pm 3.42$ $84.25\pm 3.09$ $85.96\pm 2.8$ $85.92\pm 2.75$
Adam (const), 1e-5 $4.87\pm 1.58$ $6.97\pm 2.01$ $12.38\pm 2.59$ $26.29\pm 4.03$ $56.09\pm 4.69$ $70.48\pm 4.56$ $78.41\pm 3.75$ $83.96\pm 3.11$
Table 8: Average and standard deviation evaluation accuracy for main experiment (C).
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
L3RS $72.15\pm 3.71$ $77.52\pm 3.11$ $80.22\pm 3.35$ $81.76\pm 3.27$ $83.89\pm 3.0$ $84.96\pm 2.81$ $85.95\pm 2.83$ $87.06\pm 2.79$
VeLO (ft) $66.93\pm 7.24$ $75.68\pm 3.35$ $78.26\pm 3.38$ $80.88\pm 2.98$ $83.33\pm 3.06$ $80.12\pm 12.15$ $59.12\pm 30.85$ $4.05\pm 0.37$
VeLO (og) $58.38\pm 4.31$ $54.52\pm 3.97$ $58.73\pm 4.2$ $64.48\pm 3.83$ $73.04\pm 3.49$ $78.45\pm 3.33$ $81.84\pm 3.41$ $84.44\pm 2.41$
Adam (cosine), 1e-3 $60.65\pm 4.08$ $67.66\pm 3.13$ $70.77\pm 3.74$ $74.73\pm 3.69$ $79.2\pm 3.24$ $81.41\pm 3.25$ $83.47\pm 3.13$ $85.58\pm 3.78$
Adam (cosine), 1e-4 $15.17\pm 2.9$ $38.45\pm 4.37$ $61.08\pm 3.91$ $74.22\pm 3.33$ $80.98\pm 3.45$ $83.73\pm 3.03$ $85.26\pm 3.0$ $86.8\pm 2.98$
Adam (cosine), 1e-5 $4.86\pm 1.84$ $5.88\pm 2.0$ $7.95\pm 2.25$ $13.78\pm 3.02$ $38.75\pm 4.17$ $63.86\pm 3.64$ $76.06\pm 3.29$ $82.1\pm 2.81$
Adam (const), 1e-3 $60.29\pm 4.24$ $60.9\pm 3.63$ $63.8\pm 3.57$ $68.15\pm 3.88$ $72.46\pm 3.92$ $75.25\pm 3.52$ $77.68\pm 3.61$ $80.7\pm 3.14$
Adam (const), 1e-4 $29.37\pm 4.04$ $60.46\pm 3.96$ $73.09\pm 3.48$ $78.62\pm 3.34$ $81.95\pm 3.29$ $83.27\pm 2.93$ $84.38\pm 2.99$ $85.86\pm 2.64$
Adam (const), 1e-5 $5.43\pm 1.9$ $7.8\pm 2.26$ $13.8\pm 2.96$ $30.14\pm 4.13$ $63.53\pm 3.66$ $75.85\pm 3.25$ $80.89\pm 3.24$ $84.49\pm 3.66$
Table 9: Average and standard deviation evaluation accuracy for main experiment (D).
Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
L3RS $15.77\pm 2.99$ $21.11\pm 3.52$ $25.5\pm 3.88$ $30.46\pm 4.05$ $40.32\pm 4.33$ $47.69\pm 4.26$ $54.74\pm 4.09$ $60.9\pm 4.04$
VeLO (ft) $16.18\pm 3.69$ $18.59\pm 3.27$ $18.55\pm 6.33$ $11.59\pm 8.62$ $5.49\pm 5.06$ $7.87\pm 7.24$ $4.38\pm 0.98$ $3.98\pm 0.27$
VeLO (og) $17.69\pm 3.1$ $21.04\pm 3.66$ $24.83\pm 3.72$ $30.11\pm 4.28$ $42.68\pm 4.34$ $53.73\pm 3.87$ $64.06\pm 4.14$ $73.21\pm 3.92$
Adam (cosine), 1e-3 $14.48\pm 2.49$ $19.26\pm 3.13$ $23.45\pm 3.71$ $28.14\pm 3.71$ $38.59\pm 4.42$ $46.64\pm 4.2$ $55.25\pm 4.19$ $66.46\pm 3.88$
Adam (cosine), 1e-4 $5.02\pm 1.43$ $7.46\pm 1.92$ $12.07\pm 2.49$ $17.92\pm 3.09$ $24.83\pm 3.62$ $32.58\pm 4.26$ $43.25\pm 4.41$ $56.57\pm 3.98$
Adam (cosine), 1e-5 $4.17\pm 1.18$ $4.29\pm 1.2$ $4.46\pm 1.25$ $4.97\pm 1.35$ $7.2\pm 1.96$ $12.18\pm 2.64$ $18.75\pm 3.21$ $26.31\pm 3.67$
Adam (const), 1e-3 $17.19\pm 3.05$ $20.19\pm 3.34$ $23.95\pm 3.68$ $28.74\pm 4.2$ $37.17\pm 4.62$ $44.95\pm 4.21$ $53.75\pm 4.12$ $65.26\pm 3.64$
Adam (const), 1e-4 $6.34\pm 1.72$ $12.18\pm 2.62$ $17.68\pm 2.95$ $22.36\pm 3.41$ $30.64\pm 4.12$ $40.09\pm 4.14$ $48.52\pm 4.22$ $59.35\pm 4.15$
Adam (const), 1e-5 $4.24\pm 1.23$ $4.43\pm 1.24$ $4.95\pm 1.38$ $6.37\pm 1.64$ $12.21\pm 2.6$ $18.84\pm 3.19$ $24.37\pm 3.61$ $34.44\pm 4.33$
Table 10: Average and standard deviation evaluation accuracy for main experiment (E).

### v2

@@
 L3RS is able to leverage smaller MLP networks for the per-layer operations, and can rely on base optimizers rather than generating the update step direction manually.
-These design choices result in a learned optimizer which has approximately 620×620$\times$620 × fewer parameters.
+These design choices result in a learned optimizer which will typically have two to three orders to magnitude fewer parameters compared to VeLO.
 5 Experiments
@@
 While the precise interpretation of these parameter dynamics is challenging, the parameters of L3RS are considerably more interpretable than those of most black-box learned optimizers.
-Base-optimizers Global No Embedding Per-layer MLP Embedding
-SGD only $-2.26pm0.20$ $64.48\pm 5.05$ $64.71\pm 5.00$ $68.18\pm 4.29$
-Adam only $-5.36pm0.20$ $61.44\pm 5.14$ $67.14\pm 4.57$ $68.52\pm 4.25$
-SGD, Adam $56.71\pm 4.93$ $65.70\pm 4.91$ $68.52\pm 4.40$ $\mathbf{68.93\pm 4.58}$
+Base-optimizers No Embedding Per-layer MLP Embedding
+SGD only $64.48\pm 5.05$ $64.71\pm 5.00$ $68.18\pm 4.29$
+Adam only $61.44\pm 5.14$ $67.14\pm 4.57$ $68.52\pm 4.25$
+SGD, Adam $65.70\pm 4.91$ $68.52\pm 4.40$ $\mathbf{68.93\pm 4.58}$
 Table 2: Average and standard deviation evaluation accuracy given different optimizers and per-layer strategies.
@@
 We investigate both the per-layer MLP style, as well as the single MLP without layer embeddings.
-We additionally report results for a Global method, which applies a single learning rate, $\lambda$, and direction weights, $\mu_{p}$, for all layers in the model.
 See Table 2 for results.
@@
 Improving this technique further can result in faster model training, better performance, and robust data generalization.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The main claims in the abstract are: exploring training learned optimizers for a narrow task domain, proposing a new architecture, the evaluation methods, and a brief description of baselines.
-The experiments justifying these claims are explained in the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: The scope of the domain in the paper is purposefully limited, the idea being to train on an extremely narrow task distribution.
-We do show experiments evaluating outside of this distribution, some of which perform well and others which do not.
-We propose explanations for why the method generalizes to some distributions but not others.
-We also acknowledge that while we only used Adam and SGD as directions provided to L3RS for the main training runs, we give examples and some results of using others.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: We do not include any theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: There is all information necessary to replicate the experiments in the paper.
-Code is not provided but the datasets used are public.
-The architecture proposed is sufficiently explained to reproduce without much difficulty.
-We intend to release the code after the paper’s acceptance.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24.
-Justification: There is all infomation necessary to replicate the experiments in the paper.
-Code is not provided but the datasets used are public.
-The architecture proposed is sufficiently explained to reproduce without much difficulty.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We’ve ensured to specify every hyperparameter for meta-training and task generation used to facilitate reproducibility and understanding.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: When applicable error bars (for graphs) and standard deviations (for tables) are reported.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: We describe the computational power required to perform out meta-training.
-The evaluation computation requirements are not specified, but is relatively cheap and understandable (typical Resnet-34 training requirements).
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Our paper uses only well established, public datasets.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: This method establishes a new learned optimizer architecture and type of training domain, neither of which have any obvious societal implications.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper does not release any models and only uses public datasets.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Datasets and baseline methods used are cited.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: This work does not introduce any assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
 Appendix A Checkpoint Pretraining

### v3

@@
 Every year, more and more optimization methods are proposed, making it increasingly challenging for practitioner to select the best one (Choi et al., 2019; Schmidt et al., 2021).
-Learned optimizers or learning-to-learn methods (Schmidhuber, 1987; Bengio et al., 1990, 2013) offer a potential solution to this challenge by automating the optimizer selection process.
+Learned optimizers or learning-to-learn methods (Schmidhuber, 1987; Bengio et al., 1990; 2013) offer a potential solution to this challenge by automating the optimizer selection process.
 These methods automate this process by either learning the hyperparameters of existing optimizers or developing entirely new optimization algorithms.
@@
 Meta-generalization refers to a learned optimizer’s ability to perform well on novel tasks it was not specifically trained for.
-One appraoch to achieving this is to train the optimizer on the widest possible range of tasks.
+One approach to achieving this is to train the optimizer on the widest possible range of tasks.
 Previous research focused on developing a single, adaptable algorithm by training it on a very broad domain.
@@
 By narrowing the scope of the meta-training domain, the learned optimizer can specialize and excel in the specific types of tasks it is designed for.
-Fine-tuning also simplifies the optimization problem by leveraging existing knowledge encoded in pretrained checkpoints (Wei et al., 2021) and is typically done for a relatively few iteration, since only a small alignment with the current task is needed to achieve a good results.
+Fine-tuning also simplifies the optimization problem by leveraging existing knowledge encoded in pretrained checkpoints (Wei et al., 2021) and is typically done for a relatively few iterations, since only a small alignment with the current task is needed to achieve good results.
 To achieve this, we introduce L3RS (Learned Layer-wise Learning Rate Scheduler, pronounced “lers”), a novel learned optimizer designed specifically to leverage the performance of a set of predefined base optimizers within a narrower domain.
-L3RS uses a variety of performance features, such as adaptive exponential moving averages (EMA) across different scales.
+L3RS uses a variety of performance features, such as adaptive exponential moving averages (EMA) across different time scales.
 EMAs, or model averaging, have demonstrated improved generalization in real-world applications (Tarvainen & Valpola, 2017; Izmailov et al., 2018).
@@
 This optimizer is simple and intuitive, having only a fraction of parameters compared to other black-box learned optimizers.
-  • We demonstrate that meta-training a learned optimizer on a narrow domain results in an optimizer that can outperform existing methods, with more that 50% speedup compared to another learned optimizer baseline, and more than 2×2$\times$2 × speedup compared to best performing traditional optimizers.
-  • We evaluate the meta-generalization of the proposed optimizer across several components: generalization to longer training horizon, different pretraining and model initializations, and different evaluation dataset.
+  • We demonstrate that meta-training a learned optimizer on a narrow domain results in an optimizer that can outperform existing methods, with more that $50\%$ speedup compared to another learned optimizer baseline, and more than $200\%$ speedup compared to best performing traditional optimizers.
+  • We evaluate the meta-generalization of the proposed optimizer across several components: generalization to longer training horizon, different pretraining and model initializations, and a different evaluation dataset.
 2 Preliminaries
@@
 Given a task $T:=\left\{\theta_{0},\{D^{t}_{K}\},D^{e}\right\}$, the learned optimizer $f_{\psi}$ uses optimization statistics $\Phi$ to update model parameters starting from $\theta_{0}$ for each of $\{D^{t}_{K}\}$ batches.
-After $K$ update steps, the final model parameters are evaluated on the evaluation set $D^{e}$ using meta-loss $L_{\psi}$.
+After $K$ update steps, the final model parameters are evaluated on the evaluation set $D^{e}$ using meta-loss $L_{\psi}(\theta_{K},D^{e})$.
 Right: Outer loop NES meta-training iteration.
@@
 The fitness of each candidate is then used to perform an NES update step, resulting in the next learned optimizer parameters, $\psi_{t+1}$.
-Consider a loss function $L(\theta,X)$, parameterized by $\theta$ for a given data $X$.
-Learned optimizers $f_{\psi}$, parameterized by $\psi$, provide a direction to update the model’s weights $\theta$ as
+Consider a loss function $L(\theta_{n},X)$, parameterized by weights $\theta_{n}$ for a given data $X$ at cetain optimization step $n$.
+Learned optimizers $f_{\psi}$, parameterized by $\psi$, provide a direction to update the model’s weights $\theta_{n}$ as
 $\theta_{n+1}=\theta_{n}+f_{\psi}(\Phi;\theta_{n},X),$ (1)
@@
 For example, the simplest learned optimizer might simply optimize the learning rate $\lambda$ of SGD, as $f_{\psi}(\Phi;\theta_{n},X)=\lambda\nabla L(\theta_{n},X)$, in which case $\Phi:=\{\nabla L\}$ and $\psi:=\{\lambda\}$.
-In case of Adam, the optimization parameters would be $\psi:=\{\lambda,\beta_{1},\beta_{2}\}$.
+In case of Adam, the optimization parameters would be $\psi:=\{\lambda,\beta_{1},\beta_{2}\}$, where $\lambda$ is the learning rate, $\beta_{1}$ and $\beta_{2}$ are the exponential decay rates for the first and second moments.
 In Figure 1 we show both the inner and the outer loop of a typical meta-training process.
@@
 We assume that $\theta_{0}$ is given as input, and we have no control over its generation process (unlike initialization-based meta-learning approaches, such as MAML, Finn et al., 2017).
-Learned optimizer parameters $\psi$ are evaluated by performing $K$ update steps (1) on every batch from $D^{t}_{K}$, to get $\theta_{K}$ and calculating the loss using the evaluation batch $L_{\psi}(\theta_{K},D^{e})$.
+Learned optimizer parameters $\psi$ are evaluated by performing $K$ update steps (equation 1) on every batch from $D^{t}_{K}$, to get $\theta_{K}$ and calculating the loss using the evaluation batch $L_{\psi}(\theta_{K},D^{e})$.
 We use Natural Evolution Strategies (NES, Salimans et al., 2017) to meta-train the learned optimizers.
@@
 NES is used to maximize the fitness function during meta-training.
-After the learned optimizer is trained, it can be evaluated to a new task distribution $\widetilde{T}:=\{\widetilde{\theta}_{0},\{\widetilde{D}^{t}_{K}\},\widetilde{D}^{e}\}$.
+After the learned optimizer is trained, it can be evaluated on a new task distribution $\widetilde{T}:=\{\widetilde{\theta}_{0},\{\widetilde{D}^{t}_{K}\},\widetilde{D}^{e}\}$.
 3 Motivation and Related Work
 We can separate the optimizers into two main categories.
-First category, so called black-box optimizers (Li & Malik, 2016; Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017; Sandler et al., 2021; Metz et al., 2020a, b, 2022), learns an update function $f_{\psi}$ from scratch using a custom inner optimization loop.
+First category, so called black-box optimizers (Li & Malik, 2016; Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017; Sandler et al., 2021; Metz et al., 2020a; b; 2022), learns an update function $f_{\psi}$ from scratch using a custom inner optimization loop.
 These methods often have a large number of parameters, making them prone to overfitting and stability issues (Harrison et al., 2022).
 While recent work has explored using Transformers as a meta-learner architecture (Chen et al., 2022; Moudgil et al., 2023; Jain et al., 2024), they have not yet shown a significant advantage over traditional optimizers or other learned optimizer architectures.
-A notable example of black-box optimizer is VeLO, which uses per-tensor HyperNetworks (Ha et al., 2016) represented by 512-wide LSTMs to generate parameters of per-parameter MLPs.
+A notable example of black-box optimizer is VeLO, which uses per-tensor HyperNetworks (Ha et al., 2016) represented by 512-wide LSTMs to generate parameters of per-parameter multi-layer perceptrons (MLPs).
 To compute the parameter updates, it passes the per-parameter features through the generated MLP network consisting of 2-hidden layer, 4-hidden MLP.
@@
 Such approaches include learning to adapt the hyperparameters of an underlying optimizer Daniel et al.
-(2016), or learn a schedule for the learning rate (Xu et al., 2017, 2019).
+(2016), or learn a schedule for the learning rate (Xu et al., 2017; 2019).
 In addition, these methods can include adapting the optimizer’s parameters directly (Shaban et al., 2019).
@@
 We leverage the benefits of hand-designed optimizers while incorporating the adaptability of learned components using a simple small neural network for learning the per-layer learning rate.
-Unlike existing methods that learn a fixed set of hyperparameters for a single existing optimizer, L3RS learns from a set of base-optimizers.
-Moreover, during meta-training we also search over the hyperparameters of the base-optimizer.
+Unlike existing methods that learn a fixed set of hyperparameters for a single existing optimizer, L3RS learns to utilize a set of base-optimizers.
+Moreover, during meta-training we also search over the hyperparameters of the base-optimizers.
 Almeida et al.
@@
 (2021) also propose an algorithm that combines Adam and SGD as a linear combination of their update direction.
-However, their method differs significantly in two key aspects: (1) they do not provide a mechanism for learning the mixing coefficient, instead treating it as fixed hyperparameters, and (2) their combination is applied globally across all layers, while our proposed method, L3RS, employs a layer-wise adaptive strategy.
+However, their method differs significantly in two key aspects: (1) they do not provide a mechanism for learning the mixing coefficient, instead treating it as fixed hyperparameters, and (2) their combination is applied globally across all layers, while our proposed method employs a layer-wise adaptive strategy.
 4 L3RS: Learned Layer-wise Learning Rate Scheduler
 Figure 2: Left: L3RS applied to a single layer of the target model.
-The MLP receives the norm of each direction provided, time features, target layer embedding, and EMA features of the global loss, layer gradient norm and layer parameter norm.
+The MLP receives time features, EMA features of the global loss, layer gradient norm and layer parameter norm, the target layer embedding, and the norm of each direction provided.
 The MLP outputs the weighting for each direction ($\mu_{p}$) as well as the final update norm ($\lambda$).
@@
 where $\lambda^{(l)}$ is a layer-specific update $l_{2}$-norm, $\mu_{p}^{(l)}$ are normalized mixing coefficients for the base optimizers’ directions $\sum_{p=1}^{P}\mu_{p}^{(l)}=1$, and $\hat{d}_{p}^{(l)}$ are normalized optimizers’ direction $\hat{d}_{p}^{(l)}=\frac{d_{p}^{(l)}}{\left\lVert d_{p}^{(l)}\right\rVert_{2}}$.
-Figure 2 shows the main input and output components of L3RS as well is provide a flow-chart on how the weights of the underlying model $\theta$ are updated.
+Figure 2 shows the main input and output components of L3RS as well as provides a flow-chart on how the weights of the underlying model $\theta$ are updated.
 In order to learn $\lambda^{(l)}$ and $\mu_{p}^{(l)}$, we provide L3RS with a feature vector, comprising the adaptive EMA, time, embedding features as well as the magnitude of the update directions defined below.
@@
 Adaptive Exponential Moving Averages (EMA).
-In order to capture both short-term and log-term performance trends, for each layer $l$, we maintain four EMAs with different smoothing factors $\gamma\in\{0,0.9,0.99\}$.
+In order to capture both short-term and log-term performance trends, for each layer $l$, we maintain three EMAs with different smoothing factors $\gamma\in\{0,0.9,0.99\}$.
 EMAs are updated recursively as $a^{(l)}_{i,k+1}:=\gamma_{i}a^{(l)}_{i,k}+(1-\gamma_{i})\xi^{(l)}$, where $a^{(l)}_{i,k}$ denotes the $i$th EMA for layer $l$ at time step $k$, and $\xi^{(l)}$ represents a layer-specific statistic of interest.
 Specifically, for a given layer we track the following:
-  • Log $l_{2}$-norm of weights. $\log(\left\lVert w^{(l)}\right\rVert_{2})$, where $w^{(l)}$ are the weights for layer $l$..
+  • Log $l_{2}$-norm of weights. $\log(\left\lVert w^{(l)}\right\rVert_{2})$, where $w^{(l)}$ are the weights for layer $l$.
   • Log $l_{2}$-norm of gradients. $\log(\left\lVert g^{(l)}\right\rVert_{2})$, where $g^{(l)}$ are the gradients for layer $l$.
@@
 Base-optimizer direction magnitude.
-For each base optimizer $d_{p}$, the log $l_{2}$-norm of the direction is provided: $\log(\left\lVert d_{p}\right\rVert_{2})$.
+For each base optimizer, we provide the log $l_{2}$-norm of its update: $\log(\left\lVert d_{p}\right\rVert_{2})$.
 4.2 Comparison with VeLO
-Optimizer State Memory Update Function Size
-SGD $\times0$ 61.4 KiB
-Adam 2.0×2.0$\times$2.0 × 240.0 KiB
-L3RS 2.0×2.0$\times$2.0 × 2.9 MiB
-VeLO 4.0×4.0$\times$4.0 × 418.78 MiB
-Table 1: Compute and memory costs for various optimizers.
-State memory is the ratio of the optimizer state size to the model parameter size.
-Update Function Size is the size of the compiled code to calculate the weight update given the gradients.
-Target model is ResNet-34 with 1000 classes.
+Optimizer Memory Overhead Compute Overhead
+SGD $\times0$ $0.003\%$
+Adam 2.0×2.0$\times$2.0 × $0.019\%$
+L3RS 2.0×2.0$\times$2.0 × $0.057\%$
+VeLO 4.0×4.0$\times$4.0 × $0.920\%$
+Table 1: Memory and compute overhead for various optimizers using a ResNet-34 model with 25 output classes.
+Memory Overhead is the ratio of optimizer state size to model parameter size.
+Compute overhead is the compute cost ratio of the optimizer update step to a full training step for a batch of 64 images.
+Compute cost is calculated using Jax compilation statistics.
 VeLO was designed with an efficient hypernetwork-style architecture.
@@
 5 Experiments
-To evaluate the performance of our proposed L3RS optimizer, we designed a series of image classification experiments using the ResNet-34 model.
+To evaluate the performance of our proposed L3RS optimizer, we designed a series of fine-tuning image classification experiments using the ResNet-34 model.
+Fine-Tuning.
+Our experiments focus on fine-tuning pretrained models.
+In particular, we are interested in optimization horizons on the order of hundreds or thousands of steps.
+While training can certainly extend further, this training regime is relevant in cases where the amount of training data or high computational costs can be a limiting factor.
 Model Choices.
@@
 VeLO uses a memory state of 4×4$\times$4 × the model parameters (plus additional memory from $3$ AdaFactor style accumulators).
-Table 1 shows a comparison of the memory state requirements for various optimizers, as well as the compute cost for each.
+Table 1 shows a comparison of the memory state requirements for various optimizers, as well as the compute overhead for each.
 We derive updates separately for every convolutional, dense or batch normalization layer of the model, including separate updates for kernels and biases.
@@
 With $P=2$ optimizers, L3RS outputs $333$ parameters per iteration.
-For NES training, we use a smooth exponential decay for both meta-learning rate $\alpha$ and the Gaussian noise standard deviation $\sigma$ by 0.5 every 500 generations.
+For NES meta-training, we use a smooth exponential decay for both meta-learning rate $\alpha$ and the Gaussian noise standard deviation $\sigma$ by 0.5 every 500 generations.
 We use the same antithetic sampling and fitness transformations as in Salimans et al., 2017.
@@
 Refer to caption Refer to caption
-Figure 3: Meta-evaluation of the L3RS, meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks.
+Figure 3: Meta-evaluation of L3RS meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks.
 Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right).
@@
 E. Random Initialization. Random initialization, evaluated on ImageNet25Eval dataset.
-F. Speedup of L3RS. For in-domain generalization, this shows how much faster L3RS achieves a given accuracy compared to the baselines.
+F. Speedup of L3RS in iterations. For in-domain generalization, this shows how much faster L3RS achieves a given accuracy compared to the baselines.
 We compare L3RS against the following baselines:
@@
   • Adam (const, Head). Adam optimizer with constant learning rate appled to only the final model layer (freezing the rest of the model).
-For optimizers which parameters depend on the number of training steps $K$, such as Adam (cosine), VeLO (ft) and L3RS, we perform separate evaluations across various $K$ values to assess their performance both within and beyond the meta-training regime.
-For all meta-evaluations we average the results of 100 tasks sampled from the specified task distribution. The average and standard deviation of eval accuracy across the sampled tasks is reported.
+For optimizers which depend on the number of training steps $K$, such as Adam (cosine), VeLO (ft) and L3RS, we perform separate evaluations across various $K$ values to assess their performance both within and beyond the meta-training regime.
+For all meta-evaluations we average the results of 100 tasks sampled from the specified task distribution.
+The average and standard deviation of eval accuracy across the sampled tasks is reported.
+All figures are reproduced using loss instead of accuracy in Appendix A.
 Results.
@@
 We also explore out-of-distribution performance by extending the number of evaluation steps beyond the meta-training range (indicated by the dashed line in all plots), changing the evaluation dataset to Places25Eval (B), modifying the initialization to utilize a model pretrained on Places (C), or using both the Places checkpoint and Places25Eval dataset (D).
-Figure 4: Learned mixing coefficients $\mu_{p}^{(l)}$ and $\lambda^{(l)}$ for each iteration of L3RS for ResNet-34 model.
-Top: average across all the layers.
-Bottom: individual for every learned component.
 First, comparing VeLO (ft) to VeLO (og), we observe that fine-tuning VeLO to a specific dataset and number of steps leads to a significant performance improvement compared to the general-purpose VeLO (og).
@@
 […]
-Interestingly, changing either the initialization or evaluation dataset to Places has a minimal effect on the relative perfoamnce of the optimizers, even though Places was not included in the meta-training data.
+Interestingly, changing either the initialization or evaluation dataset to Places has a minimal effect on the relative performance of the optimizers, even though Places was not included in the meta-training data.
 This highlights the robust generalization capabilities of both L3RS and the fine-tuned VeLO.
@@
 We believe this is because, while L3RS generalizes well to different checkpoints (C), the randomly initialized model may be too different and far out of distribution for L3RS to perform optimally.
+Figure 4: Visualization of learned mixing coefficients $\mu^{(l)}$ and per-layer learning rates $\lambda^{(l)}$ over 100 steps for a ResNet-34 model.
+Each layer’s type and component are distinguished by color and line type (see legend).
+The general trend shows curves moving up and to the left, indicating a transition from Adam ($\mu^{(l)}=0$) to SGD ($\mu^{(l)}=1$) and a decrease in $\lambda^{(l)}$.
+The initial step is marked with a $\bullet$ and the final step with a $\star$.
+Figure 5: Average learning rate vs. direction mix between Adam and SGD for each of the 100 steps of the L3RS optimizer.
 In Figure 3 (F) we show the speedup, in training steps, L3RS achieves over baseline methods to reach a given accuracy.
-L3RS robustly is able to demonstrate 50% speedup over VeLO (ft) and $100\%$ speedup over the best hand-designed optimizer.
-Finally, Figure 4 shows the learned parameters of L3RS during a $100$-step evaluation run.
-Interestingly, the optimizer initially heavily favors Adam direction, switching towards SGD as training progresses (with a curious sudden switch for some layers around iteration $10$).
-Moreover, some layers consistently rely on SGD or Adam direction throughout the entire run.
-While the precise interpretation of these parameter dynamics is challenging, the parameters of L3RS are considerably more interpretable than those of most black-box learned optimizers.
-Base-optimizers No Embedding Per-layer MLP Embedding
-SGD only $64.48\pm 5.05$ $64.71\pm 5.00$ $68.18\pm 4.29$
-Adam only $61.44\pm 5.14$ $67.14\pm 4.57$ $68.52\pm 4.25$
-SGD, Adam $65.70\pm 4.91$ $68.52\pm 4.40$ $\mathbf{68.93\pm 4.58}$
-Table 2: Average and standard deviation evaluation accuracy given different optimizers and per-layer strategies.
+L3RS is able to demonstrate robust 50% speedup over VeLO (ft) and $100\%$ speedup over the best hand-designed optimizer.
+The learned parameters of L3RS exhibit interesting dynamics during a 100-step evaluation (Figure 4).
+Initially, the optimizer strongly favors the Adam direction with a high learning rate for most layers.
+As training progresses, this preference shifts, transitioning towards SGD and a lower learning rate (represented by a movement towards the top-left of the plot).
+Figure 5, showing the average parameter movement across all layers, reveals several distinct phases: an initial warm-up period with an increasing learning rate, a period of relatively constant learning rate while transitioning from Adam to SGD, a phase of rapid learning rate decay, and a final convergence to the SGD direction over the last $\sim$10 steps.
+While the precise interpretation of these parameter dynamics is challenging, the L3RS parameters are considerably more interpretable than those of most black-box learned optimizers.
+Base-optimizers Embedding No Embedding Per-layer MLP Global
+SGD only $68.18\pm 4.29$ $64.48\pm 5.05$ $64.71\pm 5.00$ $64.37\pm 4.94$
+Adam only $68.52\pm 4.25$ $61.44\pm 5.14$ $67.14\pm 4.57$ $63.45\pm 4.84$
+SGD, Adam $\mathbf{68.93\pm 4.58}$ $65.70\pm 4.91$ $68.52\pm 4.40$ $66.58\pm 4.93$
+Table 2: Average and standard deviation of evaluation accuracy for different optimizers and per-layer strategies.
 EMA’s Smoothing Factors Accuracy
@@
 Ablation of Base Optimizers.
-We first show the results of using either SGD or Adam alone as the wrapped optimizer rather than both together.
+We will show the results of using either SGD or Adam alone as the wrapped optimizer rather than both together.
 Embedding Variants.
-Without providing layer embeddings to L3RS, the optimizer would have more difficulty identifying which layer it was operating on needing to rely heavily on the adaptive EMA features.
-An alternative to the layer embeddings is to meta-learn a separate MLP model for each layer of the target model.
-This strategy greatly increases the total parameter count of the optimizer.
-We investigate both the per-layer MLP style, as well as the single MLP without layer embeddings.
-See Table 2 for results.
+Without layer embeddings, L3RS struggles to distinguish between layers, relying heavily on adaptive EMA features.
+Alternatively, we can meta-learn a separate MLP for each layer of the target model.
+This significantly increases the optimizer’s parameter count.
+We compare this per-layer MLP approach with a single shared MLP (without layer embeddings).
+We additionally report results for a Global method, which uses a single learning rate ($\lambda$) and direction weights ($\mu_{p}$) for all layers.
+Results for these different optimizer configurations are presented in Table 2.
 Ablation of Adaptive EMA Input Features.
-We additionally train models with different adaptive EMA feature smoothing factors.
-While L3RS uses $\gamma\in\{0.99,0.9,0.0\}$, we also test $\gamma\in\{0.9,0.0\}$, $\gamma\in\{0.0\}$ (which corresponds to the raw features without averaging), and $\gamma\in\varnothing$ (which corresponds to not receiving the adaptive features at all).
-See Table 3 for results.
+We also investigated the impact of different smoothing factors ($\gamma$) for the adaptive exponential moving average (EMA) features.
+The standard L3RS uses $\gamma\in\{0.99,0.9,0.0\}$.
+For comparison, we trained models with $\gamma\in\{0.9,0.0\}$, $\gamma=0.0$ (representing raw features without averaging), and no adaptive EMA features ($\gamma=\emptyset$).
+The results are summarized in Table 3.
 Base Optimizer Set Accuracy
@@
 We provide results repurposing VeLO, an architecture designed for general learned optimization, as well as propose a new architecture L3RS, designed to take advantage of the narrow domain.
-We also demonstrate the robustness of this method on out of distribution tasks.
+We demonstrate the robustness of this method on out of distribution tasks.
 As fine-tuning tasks continue to become more relevant, especially for LLMs, this paradigm and architecture provides a promising direction for future research.
@@
 L3RS reaches a high fitness very rapidly then slowly improves throughout the rest of the training, likely due to the smaller number of parameters in the learned optimizer.
-Figures 7 and 7 show the example of the relative and absolute time features used by L3RS.
-Appendix C Evaluation Tables
-We provide the full evaluation tables from the main evaluation experiments Figure 5 A-E.
-Table 5 provides the results from the in-domain evaluation A.
-Table 6 provides the results for all head-only fine-tuning for evaluation A.
-Table 7 provides the results from the dataset-generalization evaluation B.
-Table 8 provides the results from the checkpoint-generalization evaluation C.
-Table 9 provides the results from the initialization and dataset-generalization evaluation D.
-Table 10 provides the results from the RandomInit evaluation E.
+Figures 8 and 8 show the example of the relative and absolute time features used by L3RS.
+Appendix C Learned mixing coefficients dynamics
+To better understand the optimization dynamics of L3RS during a 100-evaluation run, we visualize the learning rate and direction mix per layer.
+Figure 9 provides a more granular view, displaying the same relationship for each individual layer.
+Appendix D Evaluation Tables
+We provide the full evaluation tables of accuracy and loss from the main evaluation experiments Figure 5 A-E.
+We additionally reproduce Figure 5 using the evaluation losses in Figure 10.
+Table 5 and Table  6 provide the accuracy and loss results respectively from the in-domain evaluation A.
+Table 7 and Table 8 provide the accuracy and loss results respectively for all head-only fine-tuning for evaluation A.
+Table 9 and Table 10 provide the accuracy and loss results respectively from the dataset-generalization evaluation B.
+Table 11 and Table 12 provide the accuracy and loss results respectively from the checkpoint-generalization evaluation C.
+Table 13 and Table 14 provide the accuracy and loss results respectively from the initialization and dataset-generalization evaluation D.
+Table 15 and Table 16 provide the accuracy and loss results respectively from the RandomInit evaluation E.
 Figure 6: Meta-training curves for L3RS (Black), and VeLO (Blue). Left: Average population fitness. Right: Average population evaluation accuracy.
@@
 Figure 8: Absolute time features as a function of max training step (K). Each of the lines represents a time input feature to the model, generated using $K$ and $\beta_{j}$.
+Figure 9: Learned mixing coefficients $\mu_{p}^{(l)}$ and $\lambda^{(l)}$ for each iteration of L3RS for ResNet-34 model.
+Top: average across all the layers.
+Bottom: individual for every learned component.
+A $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
+Refer to caption B $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: Places25Eval
+Refer to caption C $\widetilde{\theta}_{0}$: Places,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
+Refer to caption D $\widetilde{\theta}_{0}$: Places,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: Places25Eval
+Refer to caption E $\widetilde{\theta}_{0}$: RandomInit,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
+Refer to caption F
+Refer to caption Refer to caption
+Figure 10: Meta-evaluation of the L3RS, meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks.
+Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right).
+Each marker represents model loss evaluation at that number of steps.
+Solid lines indicate the number of steps for in-distribution evaluation, while dashed lines indicate generalization to more steps than meta-training.
+A. In-domain Generalization. Both initialization and evaluation are on ImageNet.
+B. Out-of-Domain Initialization. Initialized on ImageNet, evaluated on Places25Eval.
+C. Out-of-Domain Evaluation. Initialized on Places, evaluated on ImageNet25Eval.
+D. Out-of-Domain Init & Eval. Both initialization and evaluation are on Places dataset.
+E. Random Initialization. Random initialization, evaluated on ImageNet25Eval dataset.
+F. Speedup of L3RS. For in-domain generalization, this shows how much faster L3RS achieves a given loss compared to the baselines.
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
@@
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+L3RS $1.17\pm 0.13$ $0.82\pm 0.11$ $0.67\pm 0.10$ $0.56\pm 0.10$ $0.46\pm 0.09$ $0.39\pm 0.08$ $0.38\pm 0.09$ $0.43\pm 0.11$
+VeLO (ft) $1.35\pm 0.17$ $0.90\pm 0.13$ $0.75\pm 0.11$ $0.62\pm 0.11$ $0.50\pm 0.09$ $0.45\pm 0.10$ $0.92\pm 0.96$ $3.23\pm 0.12$
+VeLO (og) $1.75\pm 0.18$ $1.81\pm 0.16$ $1.57\pm 0.14$ $1.33\pm 0.14$ $0.98\pm 0.13$ $0.74\pm 0.12$ $0.61\pm 0.11$ $0.67\pm 0.14$
+Adam (cosine), 1e-3 $1.67\pm 0.15$ $1.20\pm 0.14$ $1.01\pm 0.13$ $0.82\pm 0.13$ $0.61\pm 0.11$ $0.50\pm 0.10$ $0.45\pm 0.09$ $0.49\pm 0.13$
+Adam (cosine), 1e-4 $3.18\pm 0.06$ $2.63\pm 0.09$ $1.94\pm 0.13$ $1.23\pm 0.14$ $0.69\pm 0.11$ $0.51\pm 0.10$ $0.42\pm 0.09$ $0.40\pm 0.10$
+Adam (cosine), 1e-5 $3.56\pm 0.05$ $3.49\pm 0.05$ $3.38\pm 0.05$ $3.17\pm 0.06$ $2.60\pm 0.09$ $1.88\pm 0.13$ $1.17\pm 0.14$ $0.66\pm 0.11$
+Adam (const), 1e-3 $1.44\pm 0.15$ $1.34\pm 0.14$ $1.20\pm 0.15$ $1.00\pm 0.13$ $0.83\pm 0.12$ $0.72\pm 0.12$ $0.66\pm 0.11$ $0.70\pm 0.15$
+Adam (const), 1e-4 $2.80\pm 0.08$ $1.95\pm 0.13$ $1.24\pm 0.14$ $0.80\pm 0.12$ $0.56\pm 0.10$ $0.46\pm 0.09$ $0.43\pm 0.09$ $0.46\pm 0.12$
+Adam (const), 1e-5 $3.51\pm 0.05$ $3.38\pm 0.05$ $3.17\pm 0.06$ $2.78\pm 0.08$ $1.89\pm 0.13$ $1.17\pm 0.14$ $0.74\pm 0.11$ $0.50\pm 0.10$
+Table 6: Average and standard deviation evaluation loss for main experiment (A).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
 VeLO (og) $42.53\pm 4.73$ $67.93\pm 4.4$ $73.68\pm 3.65$ $76.35\pm 3.58$ $79.12\pm 3.45$ $80.71\pm 3.2$ $81.36\pm 3.11$ $81.6\pm 3.23$
@@
 Table 7: Average and standard deviation evaluation accuracy for all Head-only fine-tuning from main experiment (A).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+VeLO (og) $2.21\pm 0.11$ $1.11\pm 0.14$ $0.92\pm 0.14$ $0.85\pm 0.13$ $0.76\pm 0.13$ $0.68\pm 0.12$ $0.65\pm 0.12$ $0.66\pm 0.12$
+Adam (cosine), 1e-3 $3.20\pm 0.05$ $2.88\pm 0.07$ $2.45\pm 0.09$ $1.86\pm 0.11$ $1.21\pm 0.12$ $0.93\pm 0.12$ $0.77\pm 0.11$ $0.66\pm 0.11$
+Adam (cosine), 1e-4 $3.55\pm 0.05$ $3.48\pm 0.05$ $3.38\pm 0.05$ $3.23\pm 0.05$ $2.89\pm 0.06$ $2.46\pm 0.08$ $1.88\pm 0.11$ $1.21\pm 0.12$
+Adam (cosine), 1e-5 $3.60\pm 0.05$ $3.59\pm 0.05$ $3.58\pm 0.05$ $3.55\pm 0.05$ $3.48\pm 0.05$ $3.39\pm 0.05$ $3.23\pm 0.05$ $2.90\pm 0.06$
+Adam (const), 1e-3 $3.00\pm 0.06$ $2.47\pm 0.09$ $1.87\pm 0.11$ $1.35\pm 0.12$ $0.95\pm 0.12$ $0.79\pm 0.11$ $0.70\pm 0.11$ $0.64\pm 0.11$
+Adam (const), 1e-4 $3.50\pm 0.05$ $3.38\pm 0.05$ $3.23\pm 0.05$ $3.00\pm 0.06$ $2.47\pm 0.08$ $1.89\pm 0.11$ $1.35\pm 0.12$ $0.93\pm 0.12$
+Adam (const), 1e-5 $3.59\pm 0.05$ $3.58\pm 0.05$ $3.55\pm 0.05$ $3.51\pm 0.05$ $3.39\pm 0.05$ $3.24\pm 0.05$ $3.00\pm 0.06$ $2.46\pm 0.08$
+Table 8: Average and standard deviation evaluation loss for all Head-only fine-tuning from main experiment (A).
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
@@
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+L3RS $1.44\pm 0.10$ $1.02\pm 0.09$ $0.85\pm 0.09$ $0.73\pm 0.10$ $0.61\pm 0.09$ $0.55\pm 0.08$ $0.49\pm 0.08$ $0.45\pm 0.08$
+VeLO (ft) $1.73\pm 0.25$ $1.13\pm 0.10$ $0.95\pm 0.09$ $0.82\pm 0.26$ $0.66\pm 0.09$ $0.78\pm 0.51$ $3.65\pm 15.11$ $1209.28\pm 11410.45$
+VeLO (og) $1.88\pm 0.12$ $1.80\pm 0.12$ $1.57\pm 0.12$ $1.34\pm 0.10$ $1.01\pm 0.11$ $0.79\pm 0.10$ $0.64\pm 0.09$ $0.53\pm 0.08$
+Adam (cosine), 1e-3 $1.82\pm 0.10$ $1.33\pm 0.09$ $1.11\pm 0.10$ $0.91\pm 0.10$ $0.71\pm 0.09$ $0.61\pm 0.08$ $0.54\pm 0.08$ $0.47\pm 0.08$
+Adam (cosine), 1e-4 $3.19\pm 0.06$ $2.75\pm 0.06$ $2.16\pm 0.08$ $1.42\pm 0.09$ $0.85\pm 0.09$ $0.66\pm 0.09$ $0.55\pm 0.08$ $0.46\pm 0.08$
+Adam (cosine), 1e-5 $3.52\pm 0.06$ $3.46\pm 0.06$ $3.37\pm 0.06$ $3.21\pm 0.06$ $2.75\pm 0.06$ $2.10\pm 0.09$ $1.35\pm 0.10$ $0.80\pm 0.09$
+Adam (const), 1e-3 $1.61\pm 0.12$ $1.45\pm 0.11$ $1.26\pm 0.10$ $1.09\pm 0.11$ $0.90\pm 0.10$ $0.80\pm 0.10$ $0.71\pm 0.10$ $0.61\pm 0.09$
+Adam (const), 1e-4 $2.91\pm 0.06$ $2.18\pm 0.08$ $1.46\pm 0.09$ $0.98\pm 0.10$ $0.72\pm 0.09$ $0.61\pm 0.08$ $0.55\pm 0.08$ $0.48\pm 0.08$
+Adam (const), 1e-5 $3.48\pm 0.06$ $3.38\pm 0.06$ $3.21\pm 0.06$ $2.90\pm 0.06$ $2.10\pm 0.09$ $1.36\pm 0.10$ $0.90\pm 0.09$ $0.63\pm 0.09$
+Table 10: Average and standard deviation evaluation loss for main experiment (B).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
 L3RS $65.65\pm 4.61$ $73.13\pm 3.99$ $77.0\pm 4.02$ $80.33\pm 3.59$ $83.79\pm 3.06$ $85.93\pm 2.97$ $86.91\pm 2.78$ $87.04\pm 2.68$
@@
 Table 11: Average and standard deviation evaluation accuracy for main experiment (C).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+L3RS $1.34\pm 0.13$ $0.95\pm 0.12$ $0.78\pm 0.11$ $0.66\pm 0.11$ $0.54\pm 0.10$ $0.47\pm 0.09$ $0.44\pm 0.09$ $0.50\pm 0.11$
+VeLO (ft) $1.64\pm 0.27$ $1.02\pm 0.13$ $0.88\pm 0.13$ $0.73\pm 0.12$ $0.59\pm 0.10$ $0.52\pm 0.09$ $0.58\pm 0.40$ $15.24\pm 39.84$
+VeLO (og) $1.66\pm 0.16$ $1.73\pm 0.15$ $1.55\pm 0.15$ $1.31\pm 0.14$ $0.97\pm 0.14$ $0.75\pm 0.12$ $0.62\pm 0.11$ $0.71\pm 0.15$
+Adam (cosine), 1e-3 $1.75\pm 0.13$ $1.33\pm 0.15$ $1.14\pm 0.13$ $0.95\pm 0.13$ $0.72\pm 0.11$ $0.60\pm 0.10$ $0.52\pm 0.11$ $0.55\pm 0.13$
+Adam (cosine), 1e-4 $3.05\pm 0.05$ $2.63\pm 0.07$ $2.05\pm 0.10$ $1.36\pm 0.13$ $0.79\pm 0.11$ $0.59\pm 0.10$ $0.48\pm 0.09$ $0.46\pm 0.10$
+Adam (cosine), 1e-5 $3.35\pm 0.05$ $3.30\pm 0.05$ $3.22\pm 0.05$ $3.07\pm 0.05$ $2.63\pm 0.07$ $1.98\pm 0.11$ $1.28\pm 0.12$ $0.74\pm 0.11$
+Adam (const), 1e-3 $1.60\pm 0.16$ $1.48\pm 0.14$ $1.33\pm 0.15$ $1.15\pm 0.13$ $0.94\pm 0.13$ $0.81\pm 0.12$ $0.74\pm 0.12$ $0.76\pm 0.16$
+Adam (const), 1e-4 $2.79\pm 0.06$ $2.08\pm 0.10$ $1.39\pm 0.12$ $0.93\pm 0.12$ $0.65\pm 0.10$ $0.54\pm 0.09$ $0.49\pm 0.09$ $0.55\pm 0.12$
+Adam (const), 1e-5 $3.32\pm 0.05$ $3.22\pm 0.05$ $3.07\pm 0.05$ $2.77\pm 0.06$ $1.99\pm 0.11$ $1.28\pm 0.12$ $0.83\pm 0.11$ $0.56\pm 0.10$
+Table 12: Average and standard deviation evaluation loss for main experiment (C).
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
@@
 Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+L3RS $2.90\pm 0.08$ $2.72\pm 0.10$ $2.59\pm 0.10$ $2.41\pm 0.12$ $2.06\pm 0.14$ $1.79\pm 0.13$ $1.54\pm 0.14$ $1.35\pm 0.16$
+VeLO (ft) $2.93\pm 0.13$ $2.82\pm 0.09$ $2.91\pm 0.48$ $14.01\pm 28.82$ $2552.78\pm 23833.64$ $14000.60\pm 138939.16$ $20165.58\pm 184680.17$ $88.27\pm 713.91$
+VeLO (og) $2.85\pm 0.09$ $2.72\pm 0.10$ $2.60\pm 0.11$ $2.41\pm 0.13$ $1.95\pm 0.15$ $1.57\pm 0.14$ $1.22\pm 0.14$ $0.95\pm 0.15$
+Adam (cosine), 1e-3 $2.98\pm 0.06$ $2.81\pm 0.08$ $2.67\pm 0.09$ $2.50\pm 0.11$ $2.13\pm 0.14$ $1.84\pm 0.14$ $1.52\pm 0.14$ $1.13\pm 0.13$
+Adam (cosine), 1e-4 $3.26\pm 0.03$ $3.18\pm 0.03$ $3.08\pm 0.04$ $2.93\pm 0.06$ $2.68\pm 0.08$ $2.42\pm 0.11$ $2.00\pm 0.13$ $1.49\pm 0.13$
+Adam (cosine), 1e-5 $3.31\pm 0.03$ $3.30\pm 0.03$ $3.29\pm 0.03$ $3.26\pm 0.03$ $3.18\pm 0.03$ $3.07\pm 0.04$ $2.91\pm 0.06$ $2.65\pm 0.09$
+Adam (const), 1e-3 $2.88\pm 0.07$ $2.73\pm 0.09$ $2.61\pm 0.10$ $2.44\pm 0.12$ $2.14\pm 0.14$ $1.86\pm 0.15$ $1.56\pm 0.14$ $1.20\pm 0.15$
+Adam (const), 1e-4 $3.21\pm 0.03$ $3.08\pm 0.04$ $2.93\pm 0.06$ $2.76\pm 0.08$ $2.46\pm 0.11$ $2.10\pm 0.12$ $1.76\pm 0.14$ $1.37\pm 0.14$
+Adam (const), 1e-5 $3.31\pm 0.03$ $3.29\pm 0.03$ $3.26\pm 0.03$ $3.21\pm 0.03$ $3.07\pm 0.04$ $2.91\pm 0.06$ $2.72\pm 0.08$ $2.36\pm 0.11$
+Table 14: Average and standard deviation evaluation loss for main experiment (D).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
 L3RS $15.77\pm 2.99$ $21.11\pm 3.52$ $25.5\pm 3.88$ $30.46\pm 4.05$ $40.32\pm 4.33$ $47.69\pm 4.26$ $54.74\pm 4.09$ $60.9\pm 4.04$
@@
 Table 15: Average and standard deviation evaluation accuracy for main experiment (E).
+Optimizer 10-Step 25-Step 50-Step 100-Step 250-Step 500-Step 1000-Step 2500-Step
+L3RS $1.07\pm 0.09$ $0.76\pm 0.08$ $0.65\pm 0.09$ $0.58\pm 0.09$ $0.51\pm 0.08$ $0.47\pm 0.08$ $0.43\pm 0.07$ $0.40\pm 0.08$
+VeLO (ft) $1.22\pm 0.25$ $0.77\pm 0.09$ $0.69\pm 0.09$ $0.60\pm 0.09$ $0.52\pm 0.08$ $1.29\pm 3.69$ $3.53\pm 12.36$ $24563.19\pm 163714.77$
+VeLO (og) $1.38\pm 0.12$ $1.49\pm 0.12$ $1.33\pm 0.11$ $1.14\pm 0.11$ $0.85\pm 0.10$ $0.69\pm 0.09$ $0.57\pm 0.09$ $0.49\pm 0.08$
+Adam (cosine), 1e-3 $1.52\pm 0.10$ $1.11\pm 0.08$ $0.97\pm 0.10$ $0.82\pm 0.10$ $0.67\pm 0.09$ $0.58\pm 0.09$ $0.52\pm 0.08$ $0.46\pm 0.08$
+Adam (cosine), 1e-4 $2.98\pm 0.06$ $2.50\pm 0.07$ $1.85\pm 0.08$ $1.12\pm 0.09$ $0.67\pm 0.08$ $0.54\pm 0.08$ $0.47\pm 0.08$ $0.41\pm 0.07$
+Adam (cosine), 1e-5 $3.34\pm 0.06$ $3.28\pm 0.06$ $3.19\pm 0.06$ $3.01\pm 0.06$ $2.50\pm 0.07$ $1.78\pm 0.08$ $1.05\pm 0.08$ $0.63\pm 0.08$
+Adam (const), 1e-3 $1.38\pm 0.12$ $1.30\pm 0.11$ $1.17\pm 0.10$ $1.03\pm 0.11$ $0.88\pm 0.11$ $0.79\pm 0.09$ $0.70\pm 0.09$ $0.61\pm 0.10$
+Adam (const), 1e-4 $2.69\pm 0.07$ $1.89\pm 0.08$ $1.15\pm 0.09$ $0.77\pm 0.08$ $0.60\pm 0.08$ $0.53\pm 0.08$ $0.49\pm 0.08$ $0.45\pm 0.08$
+Adam (const), 1e-5 $3.30\pm 0.06$ $3.19\pm 0.06$ $3.01\pm 0.06$ $2.67\pm 0.07$ $1.78\pm 0.08$ $1.05\pm 0.08$ $0.69\pm 0.08$ $0.51\pm 0.08$
+Table 16: Average and standard deviation evaluation loss for main experiment (E).